Solve the problem with complete and proper solution. Calculate for the temperature of a mol of hydrogen gas at pressure of 2.02 x 10^4 N/m^2. A mole of hydrogen gas is 2 g. Its density is 1.5 kg/m^3. Assume that it follows the law of the ideal gas.


1
Expert's answer
2021-12-28T08:44:41-0500

1) Write down the formula of the ideal gas law:


"P=\dfrac{\rho }{\Mu}RT."


2) Express the temperature T and substitute values to find the temperature in Kelvins:


"T=\dfrac{P\Mu}{\rho R}=\dfrac{(2.02·10^4)(0.002)}{1.5 ·8.314}=3.24\text{ K}."

This is equivalent to temperature in degrees Celsius of


"t=T-273.15=-269.9\text{ K}."